OpenStudy (anonymous):

find the derivatives of y = sqrt x-4 by first principle

OpenStudy (anonymous):

\[(1/2)\sqrt{x-4}\]

OpenStudy (anonymous):

\[\left( \sqrt{x-4} \right)\prime\] This is in the form \[a\left( f(x) \right)^{b}\] Where the derivative is \[a*b*(f(x))^{b-1}*(f(x))\prime\] In our case we have, a = 1 f(x) = x-4 b = 1/2 The derivative of our function is = \[1*(1/2)*(x-4)^{(-1/2)}*1\]